# Survey on Factuality in Large Language Models: Knowledge, Retrieval and Domain-Specificity

### v1

@@ 2.2 Factuality
Both hallucinations and factuality issues in LLMs pertain to the accuracy and reliability of generated content, they address distinct aspects. Hallucinations primarily revolve around LLMs generating baseless or unwarranted content. Drawing from definitions by OpenAI (2023); Ji et al. (2023a), hallucinations can be understood as the model’s inclination to ”produce content that is nonsensical or untruthful in relation to certain sources.” This is different from factuality concerns, which emphasize the model’s ability to learn, acquire, and utilize factual knowledge. To illustrate the distinction: If an LLM, when prompted to craft ”a fairy tale about a rabbit and a wolf making friends,” produces a tale about ”a rabbit and a dog becoming friends,” it’s exhibiting hallucination. However, this isn’t necessarily a factuality error. If the generated content contains accurate information but diverges from the prompt’s specifics, it’s a hallucination but not a factuality issue. For instance, if the LLM’s output includes more details or different elements than the prompt specifies but remains factually correct, it’s a case of hallucination. Conversely, if the LLM avoids giving a direct answer, states ”I don’t know,” or provides a response that’s accurate but omits some correct details, it’s addressing factuality, not hallucination. Furthermore, it’s worth noting that hallucination can sometimes produce content that, while deviating from the original input, remains factually accurate. For a more structured comparison between the factuality issue and hallucination, refer to Table II.
Outdated information, on the other hand, focuses on instances where previously accurate information has been superseded by more recent knowledge. Lastly, domain-specificity emphasize the generation of content that requires specific, specialized knowledge. Despite these differences, all three topics contribute to our understanding of the broader factuality problem in LLMs.
@@ 2.3 Impact
A recent study (Nori et al., 2023) provides a comprehensive evaluation of GPT-4’s performance in medical competency examinations and benchmark datasets. The evaluation utilizes the text-only version of GPT-4 and investigates its ability to address medical questions without any training or fine-tuning. The assessment is conducted using the United States Medical Licensing Examination (USMLE) Kung et al. (2023) and the MultiMedQA benchmark Singhal et al. (2023), comparing GPT-4’s performance against earlier models like GPT-3.5 and models specifically fine-tuned on medical knowledge. The results demonstrate that GPT-4 significantly outperforms its predecessors, achieving scores on the USMLE that exceed the passing threshold by more than 20 points and delivering the best overall performance without specialized prompt crafting or domain-specific fine-tuning.
@@
While large language models show promise on medical datasets, the introduction of automation in the healthcare field still requires extreme caution Shen et al. (2023). Existing metrics and benchmarks are often developed for highly focused problems. Evaluating LLM outputs in supporting real-world decision-making poses challenges Singhal et al. (2023), including the stability and robustness of personalized recommendations and inferences in real-world contexts. Using large language models carries significant risks, including inaccurate ranking recommendations Hirosawa et al. (2023) (e.g., differential diagnosis) and sequencing Zhang et al. (2023c) (e.g., information gathering and testing), as well as factual errors Shen et al. (2023), particularly important omissions and erroneous responses.
3 Factuality Evaluation
@@
Evaluating the factuality of LLMs is pivotal for ensuring the reliability and trustworthiness of their generated content (Pezeshkpour, 2023; Lee et al., 2022b). As LLMs become increasingly integrated into various applications, from information retrieval to content generation, the accuracy of their outputs becomes paramount. In this section, we delve into the benchmarks used for assessing the factuality of LLMs, studies that have undertaken such evaluations, and methodologies for appraising the effectiveness of these factuality evaluators.
3.1 Factuality Evaluation Metrics
@@
We categorize these metrics into the following groups: (1) Rule-based evaluation metrics; (2) Neural evaluation metrics; (3) Hybrid evaluation metrics; and (4) Human evaluation metrics.
3.1.1 Rule-based evaluation metrics
@@
An “exact match” refers to a situation where the generated text precisely matches a specific input or reference text. This means that the LLM produces output that is identical, word-for-word, to the provided input or reference text. Exact matches are often used in NLG when you want to replicate or repeat a piece of text without any variations or alterations.
@@
Accuracy measures the overall correctness of a model’s predictions across all classes. It is the ratio of correctly predicted instances to the total number of instances.
[…]
Precision quantifies the ability of a model to correctly classify instances as positive for a specific class. It focuses on minimizing false positives. For a class $c$, the precision is given by:
[…]
Precision can be computed for each class in a multi-class problem and can be taken as the average (e.g., micro-averaging or macro-averaging) to get an overall precision score.
Recall (also known as Sensitivity or True Positive Rate) measures the ability of a model to identify all positive instances of a specific class. It focuses on minimizing false negatives. For a class $c$, the recall is given by:
[…]
Like precision, recall can be computed for each class and averaged to obtain an overall recall score.
The F-Measure is a metric that combines both precision and recall into a single score. It is especially useful for balancing precision and recall. For a class $c$, the F-Measure is calculated as:
$\mathrm{F}_{c}=\frac{2\cdot\text{ Precision }_{c}\cdot\text{ Recall }_{c}}{\text{ Precision }_{c}+\operatorname{Recall}_{c}}$ (4)
Similar to precision and recall, the F-Measure can be computed for each class and then averaged to obtain an overall F-Measure score. Common methods for averaging include micro-F1 (averaging over all instances) and macro-F1 (averaging over classes).
@@
$BS=\frac{1}{N}\sum_{i=1}^{N}(forecast_{i}-actual_{i})^{2}$ (5)
@@
(Papineni et al., 2002), also known as Bilingual Evaluation Understudy metric is commonly employed in the context of factuality evaluation. This metric calculates the co-occurrence frequency of phrases in two sentences, based on a weighted average of matched n-gram phrases. This helps in quantitatively assessing the factual consistency between the generated text and its reference. The BLEU score is computed as
$BLEU=BP*\exp\left(\frac{1}{N}*\sum_{n=1}^{N}P_{n}\right)$ (6)
where (1) $BP$ (Brevity Penalty) accounts for the length of the candidate translation. If the candidate translation is shorter than the reference, $BP$ is less than 1, which reduces the BLEU score. (2) $P_{n}$ is the n-gram precision, which measures the portion of n-grams in the candidate translation that are also present in the reference translation. (3) $N$ is the maximum order of n-grams considered (commonly up to 4 in much of the literature) and $\sum_{n=1}^{N}P_{n}$ is the sum of $\log P_{n}$ for $n$ from 1 to $N$. - $\exp$ denotes the exponential function.
@@
The Recall-Oriented Understudy for Gisting Evaluation (ROUGE) metric (Lin, 2004) serves as a measure of the similarity between the generated text and the reference text, with the similarity grounded on recall scores. Primarily used in the field of text summarization, the ROUGE metric incorporates four distinct types. These include ROUGE-n, which assesses n-gram co-occurrence statistics, and ROUGE-l, which measures the longest common subsequence. ROUGE-w provides evaluation based on the weighted longest common subsequence, while ROUGE-s measures skip-bigram co-occurrence statistics. These diverse metrics collectively provide a comprehensive measure of the factual accuracy of generated text. ROUGE score can be calculated in various ways based on the length of n-grams (unigram, bigram, etc.) used. The simplest version, ROUGE-n score can be compactly represented as:
[…]
where: (1) $RS$ denotes the set of Reference Summaries, (2) $gram_{n}$ represents an n-gram within the reference summary $S$, (3) $Count_{match}(gram_{n})$ signifies the number of times that n-gram $gram_{n}$ appears in both the generated text and the reference summary, (4) $Count(gram_{n})$ is the frequency of occurrence of the n-gram $gram_{n}$ within the reference summary.
@@
(Weller et al., 2023) is an n-gram overlap measure. It quantifies the degree to which a generated passage consists of exact spans found in a text corpus. The QUIP-Score serves to evaluate the ’grounding’ ability of LLMs, specifically assessing whether model-generated answers can be directly located within the underlying text corpus. It is defined by comparing the precision of the character n-gram from the generated output to the pre-training corpus. This is formally illustrated by generation $Y$ and the text corpus $C$:
$\operatorname{QUIP}(Y;C)=\frac{\sum_{\operatorname{gram}_{n}\in Y}\mathbb{1}_{C}\left(\operatorname{gram}_{n}\right)}{\left|\operatorname{gram}_{n}\in Y\right|},$ (8)
where $\mathbb{1}($.))))is an indicator function:
$\mathbb{1}(.)=\begin{cases}1,&\text{if }\text{gram}\in C\\ 0,&\text{otherwise}\end{cases}$ (9)
Therefore, a score of 0.5 implies that 50% of the n-grams derived from the generated text can be found in the pre-training corpus. The authors calculate a macro-average of this value over a collection of generations, which results in a single performance figure representative of a specific test dataset.
3.1.2 Neural Evaluation Metrics
@@
The automatic Dialogue Evaluation Model (ADEM) metric is a cogent tool utilized to gauge the quality of responses generated by language models in a conversation. Developed by Lowe et al. (2017), this metric trains a Hierarchical RNN, in a semi-supervised fashion, to predict ratings for the machine-generated responses. The ADEM’s assessment primarily hinges on a dialogue context, designated as $\mathbf{c}$, alongside the model’s response, labeled as $\mathbf{\hat{r}}$, and a reference response, specified as $\mathbf{r}$. These elements, when encoded via the hierarchical RNN, inform the ADEM which then predicts a score, reflecting the proximity of the model’s response to factual accuracy and relevance:
$score=(\mathbf{c}^{\top}M\mathbf{\hat{r}}+\mathbf{r}^{\top}N\mathbf{\hat{r}}-\alpha)/\beta,$ (10)
where $M,N$ are learnable parameters and $\alpha,\beta$, represent scalar constants that serve as initial values.
@@
the BERTScore metric, as introduced by Zhang et al. (2020), utilizes pre-trained contextual embeddings from BERT to gauge the similarity between two sentences. This is done by assigning contextual embeddings to a referenced sentence, ”x”, and a model-generated sentence, ”$\hat{x}$”, denoted as ”$\mathbf{x}$” and ”$\mathbf{\hat{x}}$”, respectively. The recall, precision, and F1-scores are then calculated to quantify the similarity between ”$\mathbf{x}$” and ”$\mathbf{\hat{x}}$”. This score provides a measure of the generated sentence’s factuality, and thus, the reliability of the large language model itself:
$R_{\rm BERT}$ $=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}\mathbf{x}_{i}^{\top}\mathbf{\hat{x}}_{j},$ (11)
$P_{\rm BERT}$ $=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}\mathbf{x}_{i}^{\top}\mathbf{\hat{x}}_{j},$ (12)
$F_{\rm BERT}$ $=2\frac{P_{\rm BERT}\cdot R_{\rm BERT}}{P_{\rm BERT}+R_{\rm BERT}},$ (13)
where the recall and precision elements are calculated based on a token-match approach. The recall score stems from comparing each token in the reference sentence ”$x$” to the corresponding token in the generated sentence ”$\hat{x}$”. Conversely, the precision score is derived by comparing each token in ”$\hat{x}$” to a token in ”$x$”. A greedy matching strategy is employed to pair tokens that demonstrate the highest degree of similarity. This method provides a comprehensive analysis of how precise and accurate the language model’s output aligns with the factual reference sentence.
@@
3.1.3 Human Evaluation Metrics
@@ 3.1.4 LLM-based Metrics
(Fu et al., 2023) is a new evaluation framework designed to assess the quality of output from generative AI models. To provide these evaluations, GPTScore taps into the emergent capabilities of 19 different pre-trained models, such as zero-shot instruction, and uses them to judge the generated texts. These models vary in scale from 80M to 175B. Testing across four text generation tasks, 22 aspects of evaluation, and 37 related datasets, has demonstrated that GPTScore can effectively evaluate text per instructions in natural language. This attribute allows it to sidestep challenges traditionally encountered in text evaluation, like the need for sample annotations and achieving custom, multi-faceted evaluations.
@@
TABLE III: Benchmarks for Evaluating Factuality in LLMs. This table details each benchmark’s task type, associated sub-datasets, evaluation metrics, and the performance of select representative LLMs.
Reference Task Type Dataset Metrics Performance of Representative LLMs
MMLU (Hendrycks et al., 2021)
Multi-Choice QA
Humanities,
Social,
Sciences,
STEM…
ACC
(ACC, 5-shot)
GPT-4: 86.4 GPT-3.5: 70
LLaMA2-70B: 68.9
TruthfulQA (Lin et al., 2022b)
QA
Health, Law,
Conspiracies,
Fiction…
Human Score,
GPT-judge,
ROUGE, BLEU,
MC1,MC2…
(zero-shot)
GPT-4: $\sim$29 (MC1)
GPT-3.5: $\sim$28 (MC1), 79.92(%true)
LLaMA2-70B: 53.37 (%true)
C-Eval (Huang et al., 2023b) Multi-Choice QA
STEM,
Social Science,
Humanities…
ACC
(zero-shot, average ACC)
GPT-4: 68.7 GPT-3.5: 54.4
LLaMA2-70B: 50.13
AGIEval (Huang et al., 2023b) Multi-Choice QA
Gaokao,
(geometry, Bio,
history…),
SAT, Law…
ACC
(zero-shot, average ACC)
GPT-4: 56.4 GPT-3.5: 42.9
LLaMA2-70B: 40.02
HaluEval (Li et al., 2023c)
Hallucination
Evaluation
HaluEval ACC
(general ACC)
GPT-3.5: 86.22
BigBench (Srivastava et al., 2023)
Multi-tasks (QA, NLI,
Fact Checking,
Reasoning…)
BigBench
Metric to each
type of task
(Big-Bench Hard)
GPT-3.5: 49.6
LLaMA-65B: 42.6
ALCE (Gao et al., 2023b)
Citation
Generation
ASQA, ELI5,
QAMPARI
MAUVE,
Exact Match,
ROUGE-L…
(ASQA, 3-psg, citation prec)
GPT-3.5: 73.9
LLaMA-33B: 23.0
QUIP (Weller et al., 2023) Generative QA
TriviaQA,
NQ, ELI5,
HotpotQA
QUIP-Score
Exact match
(ELI5, QUIP, null prompt)
GPT-4: 21.0
GPT-3.5: 27.7
PopQA (Mallen et al., 2023) Multi-Choice QA
PopQA,
EntityQuestions
ACC
(overall ACC)
GPT-3.5: $\sim$37.0
UniLC (Zhang et al., 2023e) Fact Checking
Climate,
Health, MGFN
ACC
F1
(zero-shot, fact tasks, average F1)
GPT-3.5: 51.62
Pinocchio (Hu et al., 2023)
Fact Checking
QA, Reasoning
Pinocchio
ACC
F1
GPT-3.5: (Zero-shot ACC: 46.8, F1:44.4)
GPT-3.5: (Few-shot ACC: 47.1, F1:45.7)
SelfAware (Yin et al., 2023) Self-evaluation
SelfAware
ACC
(instruction input, F1)
GPT-4: 75.47 GPT-3.5: 51.43
LLaMA-65B: 46.89
RealTimeQA (Kasai et al., 2022)
Multi-Choice QA,
Generative QA
RealTimeQA ACC, F1
(original setting, GCS retrieval)
GPT-3: 69.3 (ACC for MC)
GPT-3: 39.4 (F1 for generation)
FreshQA (Vu et al., 2023) Generative QA FRESHQA ACC (Human)
(strict ACC, null prompt)
GPT-4: 28.6 GPT-3.5: 26.0
3.2 Benchmarks for Factuality Evaluation
In this section, we delve into the benchmarks that are prominently employed to assess the factuality of LLMs. Specific benchmarks tailored for evaluating factuality in LLMs are tabulated in Table III.
TABLE IV: Factuality Evaluation Studies of LLMs. This table shows studies focused on factuality evaluation, and those providing valuable insights into such evaluations. The column ‘Human Eval’ denotes if the evaluation incorporated human evaluation. The ‘Granularity’ column specifies the level of evaluation, with token-level (T) and sentence-level (S) distinctions.
Reference Task Dataset Metrics Human Eval Evaluated LLMs Granularity
FActScore
(Min et al., 2023)
Biography
Generation
183 people
entities
F1
GPT-3.5
ChatGPT…
T
SelfCheckGPT
(Manakul et al., 2023)
Bio Generation WikiBio
AUC-PR,
Human Score
GPT-3,
LLaMA,
OPT,
GPT-J…
S
(Wang et al., 2023a) Open QA
NQ, TQ
ACC,
EM
GPT-3.5
ChatGPT
GPT-4
Bing Chat
S
(Pezeshkpour, 2023)
Knowledge
Probing
T-REx,
LAMA
ACC GPT3.5 T
(De Cao et al., 2021)
QA,
Fact Checking
KILT,
FEVER,
zsRE
ACC
GPT-3
FLAN-T5
S/ T
(Varshney et al., 2023)
Article
Generation
Unnamed
Dataset
ACC,
AUC
GPT3.5
Vicuna
S
FactTool
(Chern et al., 2023)
KB-based QA
RoSE
ACC
F1…
GPT-4
ChatGPT
FLAN-T5
S
(Kadavath et al., 2022) Self-evaluation
BIG Bench,
MMLU, LogiQA,
TruthfulQA,
QuALITY
TriviaQA
Lambada
ACC,
Brier Score,
RMS
Calibration
Error…
Claude T

### v3

@@ 2.2. Factuality
Therefore, we consider these three topics to fall within the scope of the factuality problem. However, it is important to note that while these topics are related, they each have a unique focus. Both hallucinations and factuality issues in LLMs pertain to the accuracy and reliability of generated content, they address distinct aspects. Hallucinations primarily revolve around LLMs generating baseless or unwarranted content. Drawing from definitions by OpenAI (2023); Ji et al. (2023a), hallucinations can be understood as the model’s inclination to ”produce content that is nonsensical or untruthful in relation to certain sources.” This is different from factuality concerns, which emphasize the model’s ability to learn, acquire, and utilize factual knowledge. To illustrate the distinction: If an LLM, when prompted to craft ”a fairy tale about a rabbit and a wolf making friends,” produces a tale about ”a rabbit and a dog becoming friends,” it’s exhibiting hallucination. However, this isn’t necessarily a factuality error. If the generated content contains accurate information but diverges from the prompt’s specifics, it’s a hallucination but not a factuality issue. For instance, if the LLM’s output includes more details or different elements than the prompt specifies but remains factually correct, it’s a case of hallucination. Conversely, if the LLM avoids giving a direct answer, states ”I don’t know,” or provides a response that’s accurate but omits some correct details, it’s addressing factuality, not hallucination. Furthermore, it’s worth noting that hallucination can sometimes produce content that, while deviating from the original input, remains factually accurate. For a more structured comparison between the factuality issue and hallucination, refer to Table 2. Outdated information, on the other hand, focuses on instances where previously accurate information has been superseded by more recent knowledge. Lastly, domain-specificity emphasize the generation of content that requires specific, specialized knowledge. Despite these differences, all three topics contribute to our understanding of the broader factuality problem in LLMs.
@@ 2.3. Impact
A recent study (Nori et al., 2023) provides a comprehensive evaluation of GPT-4’s performance in medical competency examinations and benchmark datasets. The evaluation utilizes the text-only version of GPT-4 and investigates its ability to address medical questions without any training or fine-tuning. The assessment is conducted using the United States Medical Licensing Examination (USMLE) (Kung et al., 2023) and the MultiMedQA benchmark (Singhal et al., 2023), comparing GPT-4’s performance against earlier models like GPT-3.5 and models specifically fine-tuned on medical knowledge. The results demonstrate that GPT-4 significantly outperforms its predecessors, achieving scores on the USMLE that exceed the passing threshold by more than 20 points and delivering the best overall performance without specialized prompt crafting or domain-specific fine-tuning.
@@
While large language models show promise on medical datasets, the introduction of automation in the healthcare field still requires extreme caution (Shen et al., 2023). Existing metrics and benchmarks are often developed for highly focused problems. Evaluating LLM outputs in supporting real-world decision-making poses challenges (Singhal et al., 2023), including the stability and robustness of personalized recommendations and inferences in real-world contexts. Using large language models carries significant risks, including inaccurate ranking recommendations (Hirosawa et al., 2023) (e.g., differential diagnosis) and sequencing (Zhang et al., 2023b) (e.g., information gathering and testing), as well as factual errors (Shen et al., 2023), particularly important omissions and erroneous responses.
3. Factuality Evaluation
@@
Evaluating the factuality of LLMs is pivotal for ensuring the reliability and trustworthiness of their generated content (Pezeshkpour, 2023; Lee et al., 2022b). As LLMs become increasingly integrated into various applications, from information retrieval to content generation, the accuracy of their outputs becomes paramount. In this section, we delve into the evaluation metrics and benchmarks used for assessing the factuality of LLMs, studies that have undertaken such evaluations, and domain-specific evaluation.
3.1. Factuality Evaluation Metrics
Table 3. Evaluation Metrics for the Factuality of LLMs.
Evaluation Type Metric Description, Purpose/Usage, and Reference/Findings Rule-based Exact Match Correct matches between generated and input/reference text. Common Metrics (Accuracy, Precision, Recall, F-Measure, Calibration, Brier) Accuracy: Measures overall correctness of model prediction. Precision: Ability of model to classify instances positively. Recall: Ability of model to identify all positive instances. F-Measure: Combines both precision and recall into a single score. Calibration: Agreement between predicted probabilities and observed frequencies. Brier: Measures accuracy of probabilistic predictions. MC1 and MC2 (Lin et al., 2022b) Used to identify the correct answer from multi-choice questions. BLEU (Papineni et al., 2002) Determines frequency of phrase co-occurrence in two sentences. ROUGE (Lin, 2004) Calculates similarity between reference and generated text. METEOR(Lowe et al., 2017) Comprehensive measure based on harmonic mean of unigram precision and recall. QUIP-Score (Weller et al., 2023) Calculates how much of generated text consists of exact spans found in a text corpus. Neural ADEM (Lowe et al., 2017) Uses a Hierarchical RNN to evaluate the quality of responses in language models. BERTScore (Zhang et al., 2020) Pre-trained embeddings from BERT used to evaluate sentence similarity. BLEURT (Sellam et al., 2020) Trains BERT on larger dataset and rate similarity of sentences . BARTScore (Yuan et al., 2021) Evaluates quality using pre-trained sequence-to-sequence models. Human AIS (Rashkin et al., 2023), Auto-AIS(Gao et al., 2023a) Evaluates if output is backed by evidence. FActScore (Min et al., 2023) Measures factual accuracy of LLMs breakpointing generated content into atomic facts. LLM-based GPTScore (Fu et al., 2023b) Evaluates quality of AI output. It is efficient and avoids the annotation requirement. GPT-judge (Lin et al., 2022b) Evaluates truthfulness of LLM answers. It is used in evaluating truthfulness and informativeness. Truthfulness and Informativeness (Lin et al., 2022b) Truthfulness Measures the honesty of LLM information. It is used to evaluate the factuality of information. Informativeness Evaluates the relevance and value of LLM responses LLM-Eval (Lin and Chen, 2023) Evaluates the quality of a conversation. It is adaptable in various scenarios.
@@ 3.1. Factuality Evaluation Metrics
We categorize these metrics into the following groups: (1) Rule-based evaluation metrics; (2) Neural evaluation metrics; (3) Human evaluation metrics; and (4) LLM-based evaluation metrics.
We list those metrics in Table 3.
3.1.1. Rule-based evaluation metrics
@@
An “exact match” refers to a situation where the generated text precisely matches a specific input or reference text. This means that the LLM produces output that is identical, word-for-word, to the provided input or reference text. Exact matches are often used in NLG when you want to replicate or repeat a piece of text without any variations or alterations. Exact Match is commonly used in Open-domain Question Answering (Izacard and Grave, 2021b; Wang et al., 2023h).
@@
(1) $BS=\frac{1}{N}\sum_{i=1}^{N}(forecast_{i}-actual_{i})^{2}$
@@
(Papineni et al., 2002), also known as Bilingual Evaluation Understudy metric is commonly employed in the context of factuality evaluation. This metric calculates the co-occurrence frequency of phrases in two sentences, based on a weighted average of matched n-gram phrases. This helps in quantitatively assessing the factual consistency between the generated text and its reference.
@@
The Recall-Oriented Understudy for Gisting Evaluation (ROUGE) metric (Lin, 2004) serves as a measure of the similarity between the generated text and the reference text, with the similarity grounded on recall scores. Primarily used in the field of text summarization, the ROUGE metric incorporates four distinct types. These include ROUGE-n, which assesses n-gram co-occurrence statistics, and ROUGE-l, which measures the longest common subsequence. ROUGE-w provides evaluation based on the weighted longest common subsequence, while ROUGE-s measures skip-bigram co-occurrence statistics. These diverse metrics collectively provide a comprehensive measure of the factual accuracy of generated text.
@@
(Weller et al., 2023) is an n-gram overlap measure. It quantifies the degree to which a generated passage consists of exact spans found in a text corpus. The QUIP-Score serves to evaluate the ’grounding’ ability of LLMs, specifically assessing whether model-generated answers can be directly located within the underlying text corpus. It is defined by comparing the precision of the character n-gram from the generated output to the pre-training corpus.
3.1.2. Neural Evaluation Metrics
@@
The automatic Dialogue Evaluation Model (ADEM) metric is a cogent tool utilized to gauge the quality of responses generated by language models in a conversation. Developed by Lowe et al. (2017), this metric trains a Hierarchical RNN, in a semi-supervised fashion, to predict ratings for the machine-generated responses. The ADEM’s assessment primarily hinges on a dialogue context, designated as $\mathbf{c}$, alongside the model’s response, labeled as $\mathbf{\hat{r}}$, and a reference response, specified as $\mathbf{r}$. These elements, when encoded via the hierarchical RNN, inform the ADEM which then predicts a score, reflecting the proximity of the model’s response to factual accuracy and relevance.
@@
the BERTScore metric, as introduced by Zhang et al. (2020), utilizes pre-trained embeddings from BERT to gauge the similarity between two sentences. This is done by assigning embeddings to a referenced sentence, ”x”, and a model-generated sentence, ”$\hat{x}$”, denoted as ”$\mathbf{x}$” and ”$\mathbf{\hat{x}}$”, respectively. The recall, precision, and F1-scores are then calculated to quantify the similarity between ”$\mathbf{x}$” and ”$\mathbf{\hat{x}}$”. This score provides a measure of the generated sentence’s factuality.
@@
3.1.3. Human Evaluation Metrics
@@ 3.1.4. LLM-based Metrics
(Fu et al., 2023b) is a new evaluation framework designed to assess the quality of output from generative AI models. To provide these evaluations, GPTScore taps into the emergent capabilities of 19 different pre-trained models, such as zero-shot instruction, and uses them to judge the generated texts. These models vary in scale from 80M to 175B. Testing across four text generation tasks, 22 aspects of evaluation, and 37 related datasets, has demonstrated that GPTScore can effectively evaluate text per instructions in natural language. This attribute allows it to sidestep challenges traditionally encountered in text evaluation, like the need for sample annotations and achieving custom, multi-faceted evaluations.
@@
Table 4. Benchmarks for Evaluating Factuality in LLMs. This table details each benchmark’s task type, associated sub-datasets, evaluation metrics, and the performance of select representative LLMs.
Reference Task Type Dataset Metrics Performance of Representative LLMs MMLU (Hendrycks et al., 2021) Multi-Choice QA Humanities, Social, Sciences, STEM… ACC (ACC, 5-shot) GPT-4: 86.4 GPT-3.5: 70 LLaMA2-70B: 68.9 TruthfulQA (Lin et al., 2022b) QA Health, Law, Conspiracies, Fiction… Human Score, GPT-judge, ROUGE, BLEU, MC1,MC2… (zero-shot) GPT-4: $\sim$29 (MC1) GPT-3.5: $\sim$28 (MC1), 79.92(%true) LLaMA2-70B: 53.37 (%true) C-Eval (Huang et al., 2023a) Multi-Choice QA STEM, Social Science, Humanities… ACC (zero-shot, average ACC) GPT-4: 68.7 GPT-3.5: 54.4 LLaMA2-70B: 50.13 AGIEval (Huang et al., 2023a) Multi-Choice QA Gaokao, (geometry, Bio, history…), SAT, Law… ACC (zero-shot, average ACC) GPT-4: 56.4 GPT-3.5: 42.9 LLaMA2-70B: 40.02 HaluEval (Li et al., 2023a) Hallucination Evaluation HaluEval ACC (general ACC) GPT-3.5: 86.22 BigBench (Srivastava et al., 2023) Multi-tasks (QA, NLI, Fact Checking, Reasoning…) BigBench Metric to each type of task (Big-Bench Hard) GPT-3.5: 49.6 LLaMA-65B: 42.6 ALCE (Gao et al., 2023b) Citation Generation ASQA, ELI5, QAMPARI MAUVE, Exact Match, ROUGE-L… (ASQA, 3-psg, citation prec) GPT-3.5: 73.9 LLaMA-33B: 23.0 QUIP (Weller et al., 2023) Generative QA TriviaQA, NQ, ELI5, HotpotQA QUIP-Score Exact match (ELI5, QUIP, null prompt) GPT-4: 21.0 GPT-3.5: 27.7 PopQA (Mallen et al., 2023) Multi-Choice QA PopQA, EntityQuestions ACC (overall ACC) GPT-3.5: $\sim$37.0 UniLC (Zhang et al., 2023d) Fact Checking Climate, Health, MGFN ACC F1 (zero-shot, fact tasks, average F1) GPT-3.5: 51.62 Pinocchio (Hu et al., 2023a) Fact Checking QA, Reasoning Pinocchio ACC F1 GPT-3.5: (Zero-shot ACC: 46.8, F1:44.4) GPT-3.5: (Few-shot ACC: 47.1, F1:45.7) SelfAware (Yin et al., 2023) Self-evaluation SelfAware ACC (instruction input, F1) GPT-4: 75.47 GPT-3.5: 51.43 LLaMA-65B: 46.89 RealTimeQA (Kasai et al., 2022) Multi-Choice QA, Generative QA RealTimeQA ACC, F1 (original setting, GCS retrieval) GPT-3: 69.3 (ACC for MC) GPT-3: 39.4 (F1 for generation) FreshQA (Vu et al., 2023) Generative QA FRESHQA ACC (Human) (strict ACC, null prompt) GPT-4: 28.6 GPT-3.5: 26.0
3.2. Benchmarks for Factuality Evaluation
In this section, we delve into the benchmarks that are prominently employed to assess the factuality of LLMs. Specific benchmarks tailored for evaluating factuality in LLMs are tabulated in Table 4.
Table 5. Factuality Evaluation Studies of LLMs. This table shows studies focused on factuality evaluation, and those providing valuable insights into such evaluations. The column ‘Human Eval’ denotes if the evaluation incorporated human evaluation. The ‘Granularity’ column specifies the level of evaluation, with token-level (T) and sentence-level (S) distinctions.
Reference Task Dataset Metrics Human Eval Evaluated LLMs Granularity FActScore (Min et al., 2023) Biography Generation 183 people entities F1 GPT-3.5 ChatGPT… T SelfCheckGPT (Manakul et al., 2023) Bio Generation WikiBio AUC-PR, Human Score GPT-3, LLaMA, OPT, GPT-J… S Wang et al. (2023b) Open QA NQ, TQ ACC, EM GPT-3.5 ChatGPT GPT-4 Bing Chat S Pezeshkpour (2023) Knowledge Probing T-REx, LAMA ACC GPT3.5 T De Cao et al. (2021) QA, Fact Checking KILT, FEVER, zsRE ACC GPT-3 FLAN-T5 S/ T Varshney et al. (2023) Article Generation Unnamed Dataset ACC, AUC GPT3.5 Vicuna S FactTool (Chern et al., 2023) KB-based QA RoSE ACC F1… GPT-4 ChatGPT FLAN-T5 S Kadavath et al. (2022) Self-evaluation BIG Bench, MMLU, LogiQA, TruthfulQA, QuALITY TriviaQA Lambada ACC, Brier Score, RMS Calibration Error… Claude T Chen et al. (2023b) QA NQ, WoW Factuality, Relevance, Coherence, Informativeness, Helpfulness, and Validity LLaMA, FLAN-T5, … T RAGAS (Es et al., 2023) QA WikiEval Faithfulness, Answer Relevance, Context Relevance ChatGPT T